THE C-H STRETCH SPECTRUM OF ETHYLENE IN LIQUID HELIUM DROPLETS

\underline{C. MICHAEL LINDSAY}, WILLIAM K. LEWIS, and ROGER E. MILLER, Department of Chemistry, The University of North Carolina at Chapel Hill, Chapel Hill, NC, 27599.

In this talk we report our study of ^{12}C_{2}H_{4} monomer and its clusters in ^{4}He with the new, “second generation” helium droplet/infrared spectroscopy apparatus at UNC. The region containing the ν_{11} and ν_{9} asymmetric C-H stretches of the monomer was scanned using a tunable, cw OPO aligned collinearly with the droplet beam and detected with a quadrupole mass spectrometer. Rotationally resolved spectra were observed with line widths as narrow as 0.08 cm^{−1}, FWHM. Details of the new instrument and the analysis of the spectra will be presented.